SPECTRAL DISSIMILARITIES BETWEEN AZULENE(C_{10}H_{8}) AND NAPHTHALENE(C_{10}H_{8})

\underline{MASAAKI BABA}, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto 606-8502, Japan.

Polycyclic aromatic hydrocarbons (PAHs) are of great interest in the molecular structure and excited-state dynamics, and there have been extensive spectroscopic and theoretical studies. Azulene and naphthalene are bicyclic aromatic hydrocarbons composed of odd- and even-membered rings, respectively. First, they were discriminated by a theory of mutual polarizability. ^{a} Naphthalene is an alternant hydrocarbon, but azulene is not. In contrast, spectral resemblances were found by John Platt et al., ^{b} and were explained by their simple model of molecular orbital. However, the absorption and emission feature of the S_{1} and S_{2} states is completely different each other. We have investigated each rotational and vibrational structures, and radiative and nonradiative processes by means of high-resolution spectroscopy ^{c} ^{d} and ab initio calculation. The equilibrium structures in the S_{0}, S_{1}, and S_{2} states are similar. This small structural change upon electronic excitation is common to PAH molecules composed of six-membered rings. The fluorescence quantum yield is high because radiationless transitions such as intersystem crossing (ISC) to the triplet state and internal conversion (IC) to the S_{0} state are very slow in the S_{1} state. In contrast, the S_{1} state of azulene is nonfluorescent and the S_{1} ← S_{0} excitation energy is abnormally small. We consider that the potential energy curve of a b_{2} vibration is shallower in the S_{1} state, and therefore the vibronic coupling with the S_{0} state is strong to enhance the IC process remarkably. This situation is, of course, due to its peculiar characteristics of odd-membered rings and molecular symmetry, which are completely different from the naphthalene molecule.

aC. A. Coulson and H. C. Longuet-Higgins, Proc. Roy. Soc. A, 191, 39 (1947)

bD. E. Mann, J. R. Platt, and H. B. Klevens, J. Chem. Phys., 17, 481 (1949)

cY. Semba, M. Baba, et al., J. Chem. Phys., 131, 024303 (2009)

dK. Yoshida, M. Baba, et al., J. Chem. Phys., 130, 194304 (2009)